OpenStudy (anonymous):

Consider a CD paying a 1.1% APR compounded daily. (a) Find the periodic interest rate. (b) Find the future value of the CD if you invest $3000 for a term of 5 years.

OpenStudy (tkhunny):

\(1.011^{1/365} = \)??

OpenStudy (anonymous):

uhhhm

OpenStudy (tkhunny):

It takes a calculator! Do you have one?

OpenStudy (anonymous):

Yes. Is that 1.011 times one divided by 365?

OpenStudy (wolf1728):

Is that 1.11% rate correct? (It seems pretty low).

OpenStudy (anonymous):

Yeah, that's what the question on my homework asks.

OpenStudy (anonymous):

Is anyone able to help me

OpenStudy (wolf1728):

I can help

OpenStudy (anonymous):

Im really having trouble figuring this out

OpenStudy (wolf1728):

When using an interest rate in a formula it has to be converted from a percentage so you must divide it by 100. 1.11% = .0111 annual rate = [(1+r/365)^365]-1 annual rate = [(1 +.0111/365)^365]-1 annual rate = 1.01116166291122 -1 annual rate = .01116166291122 changing it to a percentage = 1.116166291122 %

OpenStudy (anonymous):

So to find the future value I plug in numbers

OpenStudy (wolf1728):

No that's just to find the interest rate converted to annual. For calculating the total amount, you need the annuity formula. (see attached)

OpenStudy (wolf1728):

So total for 3,000 after 5 years at .01116166291122 interest Total = (3,000 * ([(1+.01116166291122)^6])/.01116166291122) - amount Total = (3000 * [(1.01116166291122)^6] / .01116166291122) - 3,000 Total = (3000 * 1.06886676311194 / .01116166291122) - 3,000 Total = (3000 * Geez it looks as if I messed up somewhere in there. Here's a page with the formulas and how to calculate and here's a page http://1728.org/annuity.htm that does all the calculation automatically. Yes, it is my site. :-)

OpenStudy (wolf1728):

Hmm I forgot the other link - here it is http://1728.org/annuity2.htm

OpenStudy (wolf1728):

Oh and the answer to that problem is 15.509.81